OpenStudy (anonymous):

Triangle DEF has vertices located at these points: Point D (-2, 5) Point E (1, 5) Point F (-2, -3) What is the area, in square units, of triangle DEF?

OpenStudy (anonymous):

area= 1/2*b*h b=sqrt[(y2-y1)^2+(x2-x1)^2] D (-2,5), E(1,5) b=sqrt[(5-5)2+(-2-1)^2]=sqrt[0+9]=3

OpenStudy (anonymous):

D(-2,5), F(-2,-3) h=sqrt[(-3-5)^2+(0)]=8

OpenStudy (anonymous):

area=1/2*3*8=12

OpenStudy (anonymous):

so twelve?